• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 10:28:21 on September21,2005

    Entry number 155493

    This entry is a followup to: 155487

    keyword=Cavities checked with spectrometer DAQ

    Cavity setpoints are ok again.  Here is the data from the spectrometer DAQ - 
    run 2714.  The cavities started working again during HAPPEX run 5256, but there 
    is some excursion in the upstream cavity - maybe just the electronics.  Anyway, 
    let's safely say the cavities started working with run 5257.
    


    FIGURE 1

    FIGURE 2